# thanks: $0$-0003-3484-911Xthanks: $0$-0002-5209-872X

### v1

@@ Abstract
at 90% confidence level (CL) from the Super-Kamiokande (SK) [4] data. Such lower bounds are generally taken to imply that any mechanism for proton decay is either at very high scales, as in Grand Unified Theories (GUTs) [5, 6, 7], or else well-sequestered from the Standard Model (SM) particles.
@@
The above considerations leads us to ask: How fast could protons have decayed in the past on Earth, or decay today far away from us? In this Letter, we attempt to answer this question. We examine a number of constraints including geological data, cosmological bounds on decaying relics, and astrophysical observations. What we find is that protons could have decayed faster, by many orders of magnitude, here on Earth until relatively recently and that this could be happening elsewhere in the Milky Way, and beyond.
We will provide simple models that show the above variations in the proton lifetime can be realized in a fairly straightforward fashion, if one invokes new ultralight scalars with feeble couplings to dark sectors. We also outline some of the expected signals of this scenario. A detection of such effects could be compelling evidence for ultralight scalars and long range forces that have yet to be discovered.
Figure 1: The various constraints on the proton lifetime. Existing constraints are in solid colors and potential future constraints are shown hatched. Each constraint probes a different region of space and time and are thus complementary. Here, SK denotes local, contemporary bounds on proton lifetime from the Super-Kamiokande experiment. The NS constraint may also improve in coming years with JWST observations.
The main idea.— We will assume that proton decay to a dark state [8, 9, 10] may be fast, if kinematically allowed. For the purposes of demonstration in this work, we will assume that the dark state is a fermion $\psi$, with no gauge charges, but having baryon number $B(\psi)=-1$ [9], which we take to be associated with a good symmetry. No cosmological abundance is assumed for $\psi$. Hence, proton decay processes such as $p\to\pi^{+}\psi$ would be baryon number preserving. On general grounds, we would then expect that the conventional decay channels containing only SM states, such as $p\to\pi^{0}e^{+}$, to be suppressed relative to the dark decay. If in the distant past, or far from the Earth, the dark fermion mass satisfies $m_{\psi}<m_{p}-m_{\pi}$, with $m_{p}\approx 0.94$ GeV and $m_{\pi}\approx 0.14$ GeV, then proton decay into a $\psi$ final state could be much faster than implied by laboratory experiments. Other decay channels, like $p\to K^{+}\psi$, can also be considered with straightforward modifications of the above.
@@
Later on, we will present models that can in principle realize cases (I) and (II) above. These models generically also lead to fast neutron decay in nuclei. Hence, one can consider the more general case of phenomenological constraints on fast nucleon decay, setting current laboratory bounds aside. Often, the decay rates for the proton and neutron are close and differ by $\mathcal{O}{(1)}$ factors. There may be environments in which neutron decays would constrain these models more tightly, such as in a neutron star (NS) with a sub-dominant proton population; in these scenarios one would also end up with an implied bound on the proton lifetime that may be stringent than a model-independent one. However, we will focus on protons in what follows, but the above proviso should be kept in mind when a particular model is considered.
@@
Geology.— Let us begin close to home, here on Earth. Stringent limits on proton decay – like those in (1) and (2) – have existed only for the past few decades. One could ask how short the proton lifetime could have been before the modern era. We focus on the heat output of the Earth and consider how much more heat it would take to melt its solid iron core.
A recent estimate using a model of heat output and transfer in the Earth suggest that the Earth’s inner core has been solid for roughly the past 1-1.3 Gyr [12]. Other estimates in the literature range from 0.7 to 4.2 Gyr [13, 14, 15]. Independent magnetic measurements suggest an increase in the Earth’s magnetic field 1-1.5 Gyr ago [16] which is consistent with this picture, although there is also a brief period of inconsistency around 0.6 Gyr ago [17].
@@
Astrophysics.— There are many potential places to look for anomalous heat generation is astrophysical settings. However, we will focus on neutron star (NS) heating as a promising signal. In cases where there is sufficient information on the NS galactic environment, observations of the effects assigned to fast proton decay can be interpreted within the context of model classes (I) and (II) above, leading to more insights on the origin of the phenomenon.
To obtain a constraint, we will use the upper bound $T<42000$ K on the temperature of the coldest known pulsar PSR J2144–3933 which is $\sim 0.3$ Gyr old, based on the observational data form the Hubble Space Telescope (HST) [19]. For this analysis, a mass of $1.4M_{\odot}$ was assumed, where $M_{\odot}\approx 2\times 10^{33}$ g is the solar mass. As we are focused on the proton lifetime, we will only consider the protons in the NS which is about $10\%$ of the total baryon content [20] across most of the star. This yields $N_{p}^{\rm NS}\sim 2\times 10^{56}$ as the number of protons in the NS. Assuming the NS is in a steady state, we expect the temperature of the NS to be given by
@@
Cosmology.— We also look to cosmology to constrain proton decay. To this end, we examine existing lower bounds on DM lifetimes. The constraints may rely on diffuse or line emission of radiation from the decay products of DM [23] or employ constraints from Cosmic Microwave Background (CMB) [24, 25]. The CMB bounds are generally stronger and also avoid issues regarding the spatial distribution of the input data. To use the DM bounds, we note that the proton can be treated as a $\sim$ GeV scale relic that carries $\Omega_{b}/\Omega_{DM}\sim 0.2$ fraction of the DM abundance. Using the results of Ref. [25], we then roughly estimate
@@
Paleo detectors.— Assuming a 10% systematic uncertainty on the kaon production rate from cosmic rays, we find that the results of Ref. [26] imply that scanning 100 g of Gyr olivine from the Earth could yield a limit on the proton lifetime of
@@
A type (I) model.— This scenario could, for example, be affected by the time evolution of a modulus $\varphi$ with a tiny mass $m_{\varphi}$ that governs the size of $m_{\psi}$. As $\varphi$ evolves, we require that $m_{\psi}(\varphi)$ be much smaller in the past, billions of years ago, allowing fast proton decay into $\psi$, but sufficiently large today $m_{\psi}(\varphi)\gtrsim m_{p}$, so that such decays are shut off. This typically implies $m_{\varphi}\gtrsim H_{0}$, where $H_{0}\approx 1.5\times 10^{-33}$ eV $\approx(14.5~{}\text{Gyr})^{-1}$ is the Hubble expansion rate today. We assume that the potential for the ultralight scalar $\varphi$ is given by
@@
A type (II) model.— As an example, take an ultralight scalar $\phi$, with a Compton wavelength $\sim$ kpc, that has a feeble coupling to dark matter (DM). Then, $\phi$ would be sourced as a background field whose value depends on DM density over galactic scales. As in case (I) above, if $m_{\psi}(\phi)<m_{p}-m_{\pi}$ then $p\to\pi^{+}\psi$ could have been fast in the past. This requires the distribution of DM near the solar system change over cosmological times, as generally expected due to local variations – see e.g. Ref. [27] – with the present value of $m_{\psi}(\phi)$ large enough to eliminate the fast proton decay channel.
Here, we adapt a model proposed in Ref. [28] for sourcing neutrino masses from the Galactic DM population. This can be straightforwardly done by introducing an interaction
@@
where $y\ll 1$ is the Yukawa coupling between the ultralight scalar $\phi$ of mass $m_{\phi}$ and the dark fermion $\psi$. As in the model of Ref. [28], we also assume that the $\phi$ couples to the main DM component $X$ through a feeble interaction
@@
with $g_{X}\ll 1$. We will assume that $m_{\phi}\sim 10^{-26}$ eV, corresponding to an $\mathcal{O}{(\rm kpc)}$ Compton wavelength, encompassing galactic distance scales. We may then deduce [28]
@@
The above implies that in the outer regions of the Milky Way, or in the inter-galactic space, where the DM density is much smaller than near the Earth, $m_{\psi}$ can be much smaller and allow fast proton decay. Similarly, if the solar system has gone through under-dense regions of the Milky Way DM halo, cosmological epoch marked by fast proton decay could have existed on Earth, in the past. For example, if such a region has a size of $\mathcal{O}{(\rm kpc)}$ – which may be typical of galactic substructures [30] – given that the solar system moves at a speed of $\sim 10^{-3}$ through the Milky Way, we may expect such an epoch to last for a few million years.
Discussion.— We summarize our results in Fig. 1 which shows the lifetimes disfavored due to different observations. As we can see, the derived bounds based on general considerations from the CMB, astrophysics, and geology yield limit that are at the level of $\sim 10^{18}$ yr, or less. A projection based on the “paleo” technique of looking for particle tracks in terrestrial minerals [26] could potentially provide a very impressive reach, not far from contemporary laboratory bounds, for stability of protons over cosmological time scales of $\mathcal{O}{(\rm Gyr)}$. However, this is only a simple projection and requires further investigation, using actual geological samples and data.
@@
We may use the results of Ref. [8] for the interaction (16), derived using a chiral perturbation formalism [32], to estimate the size of $M$ for a given proton lifetime set by $p\to\pi^{+}\psi$. If $\psi$ is relatively light, we may expect a lifetime of $\sim 6\times 10^{31}$ yr for $M=10^{15}$ GeV [8], with the lifetime scaling as $M^{4}$. Assuming a lifetime of $\sim 10^{20}$ yr, not far above the bounds derived here, we would then find $M\sim 10^{12}$ GeV, which is far below the typical GUT scales $\gtrsim 10^{15}$ GeV.
@@
Summary.— In this Letter, we have taken the point of view that the proton decay rate may vary throughout space and time, due to its dependence on an ultralight modulus. This modulus could possibly be undergoing cosmological evolution, resulting in a time varying effect. However, this field may also be sourced, for example by dark matter, in which case one could generally expect both spatial and temporal modulations of the proton decay rate. In this type of scenario, the lower bound on the proton lifetime deduced from modern terrestrial data could be a transient effect, and that protons could be decaying much faster far from the Earth or at much earlier times. We then inquired what lower bounds on the proton lifetime could be derived, independently of laboratory constraints.

### v2

@@ I Introduction
at 90% confidence level (CL) from the Super-Kamiokande (SK) [4] data. Such lower bounds are generally taken to imply that any mechanism for proton decay is either at very high scales, as in Grand Unified Theories (GUTs) [5, 6, 7], or else well sequestered from the Standard Model (SM) particles.
@@
The above considerations lead us to ask: How fast could protons have decayed in the past on Earth, or decay today far away from us? In this paper, we attempt to answer this question. We examine a number of constraints including geological data, cosmological bounds on decaying relics, and astrophysical observations. What we find is that protons could have decayed faster, by many orders of magnitude, here on Earth until relatively recently and that this could be happening elsewhere in the Milky Way, and beyond.
We will provide simple models that show the above variations in the proton lifetime can be realized in a fairly straightforward fashion, if one invokes new ultralight scalars with feeble couplings to dark sectors. Ultralight bosons have rich phenomenology touching on dark matter – see Ref. [8] for a recent review on ultralight dark matter – and lead to unique phenomenological signatures such as superradiance, see Ref. [9] for a comprehensive review of superradiance. We also outline some of the expected signals of proton decay due to a coupling to an ultralight scalar. A detection of such effects could be compelling evidence for ultralight scalars and long range forces that have yet to be discovered.
Figure 1: The various constraints on the proton lifetime along with the approximate timescale; see the text for details. Existing constraints are in solid colors and potential future constraints are shown hatched. Each constraint probes a different region of space and time and are thus complementary. The NS constraint may also improve in coming years with JWST observations.
II The main idea
We will assume that proton decay to a dark state [10, 11, 12] may be fast, if kinematically allowed. For the purposes of demonstration in this work, we will assume that the dark state is a fermion $\psi$, with no gauge charges, but having baryon number $B(\psi)=+1$ [11], which we take to be associated with a good symmetry. No cosmological abundance is assumed for $\psi$, and for parameters relevant to the constraints derived below, we find this assumption is valid by many orders of magnitude. Hence, proton decay processes such as $p\to\pi^{+}\psi$ would be baryon number preserving. On general grounds, we would then expect that the conventional decay channels containing only SM states, such as $p\to\pi^{0}e^{+}$, to be suppressed relative to the dark decay. If in the distant past, or far from the Earth, the dark fermion mass satisfies $m_{\psi}<m_{p}-m_{\pi}$, with $m_{p}\approx 0.94$ GeV and $m_{\pi}\approx 0.14$ GeV, then proton decay into a $\psi$ final state could be much faster than implied by laboratory experiments. Other decay channels, like $p\to K^{+}\psi$, can also be considered with straightforward modifications of the above.
@@ II The main idea
Later on, we will present models that can in principle realize cases (I) and (II) above. These models generically also lead to fast neutron decay in nuclei. Hence, one can consider the more general case of phenomenological constraints on fast nucleon decay, setting current laboratory bounds aside. Often, the decay rates for the proton and neutron are close and differ by $\mathcal{O}{(1)}$ factors. There may be environments in which neutron decays would constrain these models more tightly, such as in a neutron star (NS) with a sub-dominant proton population; in these scenarios one would also end up with an implied bound on the proton lifetime that may be more stringent than a model-independent one. However, we will focus on protons in what follows, but the above proviso should be kept in mind when a particular model is considered.
III Novel constraints
@@
III.1 Geology
Let us begin close to home, here on Earth. Stringent limits on proton decay – like those in (1) and (2) – have existed only for the past few decades. One could ask how short the proton lifetime could have been before the modern era. We focus on the heat output of the Earth and consider how much more heat it would take to melt its solid iron core.
A recent estimate using a model of heat output and transfer in the Earth suggest that the Earth’s inner core has been solid for roughly the past 1-1.3 Gyr [14]. Other estimates in the literature range from 0.7 to 4.2 Gyr [15, 16, 17]. Independent magnetic measurements suggest an increase in the Earth’s magnetic field 1-1.5 Gyr ago [18], which is consistent with this picture, although there is also a brief period of inconsistency around 0.6 Gyr ago [19].
@@ III.1 Geology
III.2 Astrophysics
There are many potential places to look for anomalous heat generation is astrophysical settings. However, we will focus on neutron star (NS) heating as a promising signal. In cases where there is sufficient information on the NS galactic environment, observations of the effects assigned to fast proton decay can be interpreted within the context of model classes (I) and (II) above, leading to more insights on the origin of the phenomenon.
To obtain a constraint, we will use the upper bound $T<42000$ K on the temperature of the coldest known pulsar PSR J2144–3933 which is $\sim 0.3$ Gyr old, based on the observational data form the Hubble Space Telescope [21]. For this analysis, a mass of $1.4M_{\odot}$ was assumed, where $M_{\odot}\approx 2\times 10^{33}$ g is the solar mass. As we are focused on the proton lifetime, we will only consider the protons in the NS which is about $10\%$ of the total baryon content [22] across most of the star. This yields $N_{p}^{\rm NS}\sim 2\times 10^{56}$ as the number of protons in the NS. Assuming the NS is in a steady state, we expect the temperature of the NS to be given by
@@ III.2 Astrophysics
III.3 Cosmology
We also look to cosmology to constrain proton decay. To this end, we examine existing lower bounds on DM lifetimes. The constraints may rely on diffuse or line emission of radiation from the decay products of DM [25] or employ constraints from Cosmic Microwave Background (CMB) [26, 27]. The CMB bounds are generally stronger and also avoid issues regarding the spatial distribution of the input data. To use the DM bounds, we note that the proton can be treated as a $\sim$ GeV scale relic that carries $\Omega_{b}/\Omega_{DM}\sim 0.2$ fraction of the DM abundance. Using the results of Ref. [27], we then roughly estimate
@@ III.3 Cosmology
We briefly mention that Big Bang Nucleosynthesis (BBN) is also sensitive to fast proton decay at times even before the CMB. Due to lower precision measurements and shorter timescales in play, however, the constraints are many orders of magnitude weaker than those of the CMB. In addition, observations of Pop-II stars in the Milky Way, which formed very early in the Universe ($\lesssim 1$ Gyr after the Big Bang) and have low metalicity [28, 29], can constrain the proton lifetime at the $\tau_{p}\sim 10^{15}$ yr level. This is due to the inferred abundance of boron in very old stars [28], which can be enhanced by the decay of carbon, as one of its protons decays. The above bound is less competitive than other constraints in the Milky Way, but does probe the intermediate ages of the Universe, before the geological time scales considered here.
III.4 Paleo detectors
Assuming a 10% systematic uncertainty on the kaon production rate from cosmic rays, we find that the results of Ref. [30] imply that scanning 100 g of Gyr olivine from the Earth could yield a limit on the proton lifetime of
@@ III.4 Paleo detectors
Another terrestrial constraint comes from the extremely low ratio of 14C/12C $\sim 2\times 10^{-18}$ abundance in natural gas measured by BOREXINO [31]. Proton decay would feed an additional source of 14C, which itself has a half life of 5700 years, from 15N. The 15N abundance in the Earth is within $1\%$ of that in the atmosphere and is $4\times 10^{-3}$ [32]. The carbon to nitrogen ratio in the Earth varies [33], but we conservatively take it to be $\sim 1$. Combined, we find that this places a limit on the lifetime of a proton at the level of $10^{19}$ years, but is only applicable over timescales of the last few thousand years. That is, if fast proton decay stopped much earlier than that, as may be expected in typical underlying models, the produced 14C abundance would have been efficiently depleted below the observed level.
There are additional means of probing anomalously fast proton decay such as looking at decay products from Andromeda or dwarf spheroidals [34, 35]; we find that these constraints are not competitive with those listed here, but could conceivably be relevant in alternative implementations of our idea.
IV Two models
@@
IV.1 A type (I) model
This scenario could, for example, be effected by the time evolution of a modulus $\varphi$ with a tiny mass $m_{\varphi}$ that governs the size of $m_{\psi}$. As $\varphi$ evolves, we require that $m_{\psi}(\varphi)$ be much smaller in the past, billions of years ago, allowing fast proton decay into $\psi$, but sufficiently large today $m_{\psi}(\varphi)\gtrsim m_{p}$, so that such decays are shut off. This typically implies $m_{\varphi}\gtrsim H_{0}$, where $H_{0}\approx 1.5\times 10^{-33}$ eV $\approx(14.5~{}\text{Gyr})^{-1}$ is the Hubble expansion rate today. We assume that the potential for the ultralight scalar $\varphi$ is given by
@@ IV.1 A type (I) model
IV.2 A type (II) model
As an example, take an ultralight scalar $\phi$, with a Compton wavelength $\sim$ kpc, that has a feeble coupling to dark matter (DM). Then, $\phi$ would be sourced as a background field whose value depends on DM density over galactic scales. As in case (I) above, if $m_{\psi}(\phi)<m_{p}-m_{\pi}$ then $p\to\pi^{+}\psi$ could have been fast in the past. This requires the distribution of DM near the solar system change over cosmological times, as generally expected due to local variations – see e.g. Ref. [36] – with the present value of $m_{\psi}(\phi)$ large enough to eliminate the fast proton decay channel.
Here, we adapt a model proposed in Ref. [37] for sourcing neutrino masses from the Galactic DM population, but we apply it to generating a spatially varying mass for $\psi$ sourced by the DM, $X$, via a long-range mediator, $\phi$. This can be straightforwardly done by introducing an interaction
@@ IV.2 A type (II) model
where $y\ll 1$ is the Yukawa coupling between the ultralight scalar $\phi$ of mass $m_{\phi}$ and the dark fermion $\psi$. As in the model of Ref. [37], we also assume that the $\phi$ couples to the DM, $X$, through a feeble interaction
@@
with $g_{X}\ll 1$. We will assume that $m_{\phi}\sim 10^{-26}$ eV, corresponding to an $\mathcal{O}{(\rm kpc)}$ Compton wavelength, encompassing galactic distance scales. One can show that the induced $\phi$ background is given by
$\phi\sim\frac{g_{X}n_{X}}{m_{\phi}^{2}}\,,$ (15)
where $n_{X}$ is the DM number density. Substituting the above in Eq. (13), we may then deduce [37]
@@
The above implies that in the outer regions of the Milky Way, or in the intergalactic space, where the DM density is much smaller than near the Earth, $m_{\psi}$ can be much smaller and allow fast proton decay. Similarly, if the solar system has gone through underdense regions of the Milky Way DM halo, cosmological epochs marked by fast proton decay could have existed on Earth, in the past. For example, if such a region has a size of $\mathcal{O}{(\rm kpc)}$ – which may be typical of galactic substructures [39] – given that the solar system moves at a speed of $\sim 10^{-3}$ through the Milky Way, we may expect such an epoch to last for a few million years.
V Discussion
We summarize our results in Fig. 1 which shows the lifetimes disfavored due to different observations. As we can see, the derived bounds based on general considerations from the CMB, astrophysics, and geology yield limits that are at the level of $\sim 10^{18}$ yr, or less. A projection based on the “paleo” technique of looking for particle tracks in terrestrial minerals [30] could potentially provide a very impressive reach, not far from contemporary laboratory bounds, for stability of protons over cosmological time scales of $\mathcal{O}{(\rm Gyr)}$. However, this is only a simple projection and requires further investigation, using actual geological samples and data.
@@ V Discussion
We may use the results of Ref. [10] for the interaction (17), derived using a chiral perturbation formalism [41], to estimate the size of $M$ for a given proton lifetime set by $p\to\pi^{+}\psi$. If $\psi$ is relatively light, we may expect a lifetime of $\sim 6\times 10^{31}$ yr for $M=10^{15}$ GeV [10], with the lifetime scaling as $M^{4}$. Assuming a lifetime of $\sim 10^{20}$ yr, not far above the bounds derived here, we would then find $M\sim 10^{12}$ GeV, which is far below the typical GUT scales $\gtrsim 10^{15}$ GeV. We also note that in a GUT model, a time or spatial variation in the proton decay could also correlate the time variation of other fundamental parameters such as $\alpha$, see e.g. [42, 43, 44, 45, 46, 47, 48, 49, 50].
@@
Finally, we mention that it is conceivable to have a positive signal for time or spatial dependent proton decay. One approach is a correlated set of measurements such as a geological period of anomalous heating of the Earth’s core combined with paleo measurements from different epochs that show no proton decay up to a certain period and do show proton decay over the same period as the Earth’s core heating at a consistent rate. Another approach is if anomalous neutron star heating is observed that is only present in neutron stars in low dark matter densities which is a trend that goes in the opposite direction as typical dark matter heating models. Other possible combinations of data sets may also lead to convincing positive signals.
VI Summary
In this paper, we have taken the point of view that the proton decay rate may vary throughout space and time, due to its dependence on an ultralight modulus. This modulus could possibly be undergoing cosmological evolution, resulting in a time varying effect. However, this field may also be sourced, for example by dark matter, in which case one could generally expect both spatial and temporal modulations of the proton decay rate. In this type of scenario, the lower bound on the proton lifetime deduced from modern terrestrial data could be a transient effect, and that protons could be decaying much faster far from the Earth or at much earlier times. We then inquired what lower bounds on the proton lifetime could be derived, independently of laboratory constraints.